# Which gravitational lensing degeneracies are broken in wave-optics?

### v1

@@ Abstract
This paper studies gravitational lensing degeneracies in the wave-optics regime, focusing on lensed gravitational waves (GWs). Considering lensing degeneracies as re-scaling (or transformations) of arrival time delay surface, we can divide them into local and global types. Local degeneracies only affect the time delay surface in localized regions, whereas global degeneracies re-scale the whole time delay surface by a constant while keeping the various observed image properties unchanged. We show that local degeneracies can be broken in the wave-optics regime since lensing effects become sensitive to the overall arrival time delay surface and not only to the time delay values at the image positions. On the other hand, global degeneracies (such as similarity transformation and mass-sheet degeneracy) multiply the amplification factor by a constant factor (let us say, $\lambda$). However, in GW lensing, as the GW signal amplitude depends on the source distance, it turns out that $\lambda$ is completely degenerate with the Hubble constant, similar to what we see in geometric optics. Hence, with the lensing of GWs, global degeneracies are as hard to break in wave optics as they are in geometric optics.
@@ 1 Introduction
Studying GW lensing in the wave-optics regime gives us access to additional information compared to the lensing of EM waves in the geometric optics regime. First, lensing effects (such as amplification factor; Takahashi & Nakamura, 2003) become sensitive to the frequency of the GW signal. Second, we observe the amplitude of the GW signal, meaning that we also have access to the phase information, which is lost in the lensing of EM waves where we observe flux (=|amplitude|2). Owing to these additional pieces of information at our disposal, we can ask: what more can we learn about the lens while studying lensing in the wave-optics regime? Can it especially help us break various lensing degeneracies originally encountered in the lensing of EM waves in geometric optics regimes? Degenracies in gravitational lensing refer to certain transformations that leave various observed image properties unchanged (Falco et al., 1985; Gorenstein et al., 1988). These transformations can be thought of as re-scaling (or transformations) of the arrival time delay surface and can be divided into global and local types (Saha, 2000). Global degeneracies re-scale the arrival time delay surface by a constant factor such as similarity and mass-sheet transformations (e.g., Wagner, 2018). On the other hand, local degeneracies affect the arrival time surface in localized regions such that all lensing observables (in geometric optics regime) are invariant under these transformations, such as monopole degeneracy (e.g., Liesenborgs et al., 2008; Liesenborgs & De Rijcke, 2012). In this work, we look at the effects of lensing degeneracies on the lensed GW signals to determine which one of these can be broken under wave-optics approximation.
@@ 2 Lensing basics in wave-optics regime
$ds^{2}=-(1+2U)c^{2}dt^{2}+a^{2}(1-2U)d\boldsymbol{r}^{2},$ (1)
@@
$(c^{2}\nabla^{2}+\omega^{2})\tilde{\phi}=4\omega^{2}U\tilde{\phi},$ (2)
@@ 3 Solving $F(f)$
where the summation is over all contours of $\tau^{\prime}$. Also, we can see that $F(\tau^{\prime})$ is a smooth function except for critical values of $\tau_{d}(\boldsymbol{x})$, i.e., time delay values corresponding to lensed images in geometric optics. We refer readers to Ulmer & Goodman (1995) on how to deal with these singularities. Another point is that we can only calculate $F(t)$ for a finite range of values whereas the integral in Equation (11) runs in range [,+][$\infty$,$\infty$][ - ∞ , + ∞ ]. To deal with this, we use a cosine apodization function as discussed in Diego et al. (2019) and Mishra et al. (2021). The above method is validated and used multiple previous works (Mishra et al., 2021; Meena et al., 2022; Meena, 2024).
@@
As discussed in Saha (2000), lensing degeneracies can be understood as the transformation of the arrival time delay surface, $t_{d}(\boldsymbol{x},\boldsymbol{y})$ given by Equation (4) such that the various observed image properties remain unchanged. These transformations can be broadly divided into global and local transformations. Global transformation scales the arrival time delay as a whole, whereas local transformation only modifies a part of it. Examples of global degeneracies are similarity transformation and mass-sheet degeneracy (MSD; e.g., Gorenstein et al., 1988). Similarity transformation only scales the arrival time delay surface by a constant factor (let us say, $\lambda$), leaving image positions ($\boldsymbol{x}_{i}$) and magnifications ($\mu_{i}$) unchanged, i.e., $(t_{d,\lambda},\boldsymbol{x}_{i,\lambda},\mu_{i,\lambda})=(\lambda t_{d},\boldsymbol{x}_{i},\mu_{i})$. MSD, on the other hand, also scales the magnification factors such that their ratios are still the same, i.e., $(t_{d,\lambda},\boldsymbol{x}_{i,\lambda},\mu_{i,\lambda})=(\lambda t_{d},\boldsymbol{x}_{i},\mu_{i}/\lambda^{2})$. Examples of local degeneracies include monopole degeneracy and generalized mass-sheet degeneracy (gMSD) [^1] Here, we call gMSD a local degeneracy since we need to locally modify the time delay surface by adding monopoles. But it also does an overall rescaling of surface density similar to MSD., or any other transformation concocted to modify the arrival time delay surface such that they leave the observable image properties unchanged. In this section, we look at three degeneracies, namely, monopole, MSD, and gMSD, in the wave-optics regime to determine which one of these are broken.
@@ 4.1 Monopole degeneracy
The monopole degeneracy was first introduced in Saha (2000) and was further discussed in detail in Liesenborgs et al. (2008); Liesenborgs & De Rijcke (2012). The monopole degeneracy allows us to alter the lens mass distribution in an axis-symmetric manner in regions of the image plane which do not contain any images such that the observed image properties remain unchanged.
@@ 4.3 Generalized mass-sheet degeneracy (gMSD)
As discussed in the previous sub-section, in MSD, we add a constant mass density, which is a multiple of the critical density ($\Sigma_{\rm cr}$). However, exactly due to the fact that it is multiple of critical density, it can be broken if we have two lensed sources since critical density is redshift dependent. Liesenborgs & De Rijcke (2012) showed that one could construct a so-called generalized mass-sheet degeneracy (gMSD), which cannot be broken even in the presence of multiple lensed sources. This is done by first rescaling $\kappa(\boldsymbol{x})$ for each lensed source mass-sheet transformation and then adding monopoles to compensate for each mass-sheet such that each of the lensed sources will only be subjected to their own MSDs. However, as shown in sub-section 4.1, adding monopoles essentially modifies the time delay surface (although not at the image position), and these changes will again be reflected in the $F(t)$ and $F(f)$ curves. Hence, we can again argue that, in wave-optics, the gMSD will be broken.
@@ 5 Summary & Conclusion
In this work, we have studied various lensing degeneracies in the wave-optics regime, primarily focusing on lensed GW signals. We have explicitly shown that monopole degeneracy breaks in the wave-optics regime since its effects are captured in the amplification factor. Based on that, we argue that any local degeneracy can be broken in the wave-optics regime since, by definition, they will modify the arrival time delay surface (although not at the image positions) as well as $F(f)$.
@@ Acknowledgements
I am grateful to Prasenjit Saha for many useful discussions and comments. I also thank Jose Diego for useful comments. I acknowledge support by grant 2020750 from the United States-Israel Binational Science Foundation (BSF) and grant 2109066 from the United States National Science Foundation (NSF), and by the Ministry of Science & Technology, Israel. This research has made use of NASA’s Astrophysics Data System Bibliographic Services.

### v2

@@ Abstract
This paper studies gravitational lensing degeneracies in the wave-optics regime, focusing on lensed gravitational waves (GWs). Considering lensing degeneracies as re-scaling (or transformations) of arrival time delay surface, we can divide them into local and global types. Local degeneracies only affect the time delay surface in localized regions, whereas global degeneracies re-scale the whole time delay surface by a constant while keeping the various observed image properties unchanged. We show that local degeneracies can be broken in the wave-optics regime since lensing effects become sensitive to the overall arrival time delay surface and not only to the time delay values at the image positions. On the other hand, global degeneracies (such as similarity transformation, prismatic transformation, and mass-sheet degeneracy) multiply the amplification factor by a constant factor (let us say, $\lambda$). However, in GW lensing, as the GW signal amplitude depends on the source distance, it turns out that $\lambda$ is completely degenerate with the Hubble constant, similar to what we see in geometric optics. Hence, with the lensing of GWs, global degeneracies are as hard to break in wave optics as they are in geometric optics.
@@ 1 Introduction
Studying GW lensing in the wave-optics regime gives us access to additional information compared to the lensing of EM waves in the geometric optics regime. First, lensing effects (such as amplification factor; Takahashi & Nakamura, 2003) become sensitive to the frequency of the GW signal. Second, we observe the amplitude of the GW signal, meaning that we also have access to the phase information, which is lost in the lensing of EM waves where we observe flux (=|amplitude|2). That said, wave effects may also show for coherent EM point sources such as fast radio bursts or pulsars (e.g., Jow et al., 2020) or in plasma lensing (e.g., Jow et al., 2023). Owing to these additional pieces of information at our disposal, we can ask: what more can we learn about the lens while studying lensing in the wave-optics regime? Can it especially help us break various lensing degeneracies originally encountered in the lensing of EM waves in geometric optics regimes? Degeneracies in gravitational lensing refer to certain transformations that leave various observed image properties unchanged (Falco et al. 1985; Gorenstein et al. 1988; also see Figure 4 in Wagner 2018). These transformations can be thought of as re-scaling (or transformations) of the arrival time delay surface and can be divided into global and local types (Saha, 2000). Global degeneracies re-scale the arrival time delay surface by a constant factor such as similarity, prismatic, and mass-sheet transformations (e.g., Wagner, 2018). On the other hand, local degeneracies affect the arrival time surface in localized regions such that all lensing observables (in geometric optics regime) are invariant under these transformations, such as monopole degeneracy (e.g., Liesenborgs et al., 2008a; Liesenborgs & De Rijcke, 2012). In this work, we look at the effects of lensing degeneracies on the lensed GW signals to determine which one of these can be broken under wave-optics approximation.
@@ 2 Lensing basics in wave-optics regime
$ds^{2}=-\left(1+\frac{2U}{c^{2}}\right)c^{2}dt^{2}+a^{2}\left(1-\frac{2U}{c^{2}}\right)d\boldsymbol{r}^{2},$ (1)
@@
$(c^{2}\nabla^{2}+\omega^{2})\tilde{\phi}=4\omega^{2}c^{-2}U\tilde{\phi},$ (2)
@@ 3 Solving $F(f)$
where the summation is over all contours of $\tau^{\prime}$. Also, we can see that $F(\tau^{\prime})$ is a smooth function except for critical values of $\tau_{d}(\boldsymbol{x})$, i.e., time delay values corresponding to lensed images in geometric optics. We refer readers to Ulmer & Goodman (1995) on how to deal with these singularities. Another point is that we can only calculate $F(t)$ for a finite range of values whereas the integral in Equation (11) runs in range [,+][$\infty$,$\infty$][ - ∞ , + ∞ ]. To deal with this, we use a cosine apodization function as discussed in Diego et al. (2019) and Mishra et al. (2021). The above method is validated and used multiple previous works (Mishra et al., 2021; Meena et al., 2022; Meena, 2024). Other methods to solve the above integral are discussed in (e.g., Diego et al., 2019; Villarrubia-Rojo et al., 2024).
@@
As discussed in Saha (2000), lensing degeneracies can be understood as the transformation of the arrival time delay surface, $t_{d}(\boldsymbol{x},\boldsymbol{y})$ given by Equation (4) such that the various observed image properties remain unchanged. These transformations can be broadly divided into global and local transformations. Global transformation scales the arrival time delay as a whole, whereas local transformation only modifies a part of it. Examples of global degeneracies are similarity transformation, prismatic transformation and mass-sheet degeneracy (MSD; e.g., Gorenstein et al., 1988). Similarity transformation only scales the arrival time delay surface by a constant factor (let us say, $\lambda$), leaving image positions ($\boldsymbol{x}_{i}$) and magnifications ($\mu_{i}$) unchanged, i.e., $(t_{d,\lambda},\boldsymbol{x}_{i,\lambda},\mu_{i,\lambda})=(\lambda t_{d},\boldsymbol{x}_{i},\mu_{i})$. Prismatic transformation leaves all of the above three observables unchanged, i.e., $(t_{d,\lambda},\boldsymbol{x}_{i,\lambda},\mu_{i,\lambda})=(t_{d},\boldsymbol{x}_{i},\mu_{i})$. MSD, in addition to time delay, also scales the magnification factors such that their ratios are still the same, i.e., $(t_{d,\lambda},\boldsymbol{x}_{i,\lambda},\mu_{i,\lambda})=(\lambda t_{d},\boldsymbol{x}_{i},\mu_{i}/\lambda^{2})$. Examples of local degeneracies include monopole degeneracy and generalized mass-sheet degeneracy (gMSD) [^1] Here, we call gMSD a local degeneracy since we need to locally modify the time delay surface by adding monopoles. But it also does an overall rescaling of surface density similar to MSD., or any other transformation concocted to modify the arrival time delay surface such that they leave the observable image properties unchanged. In this section, we look at three degeneracies, namely, monopole, MSD, and gMSD, in the wave-optics regime to determine which one of these are broken.
@@ 4.1 Monopole degeneracy
The monopole degeneracy was first introduced in Saha (2000) and was further discussed in detail in Liesenborgs et al. (2008a); Liesenborgs & De Rijcke (2012). The monopole degeneracy allows us to alter the lens mass distribution in an axis-symmetric manner in regions of the image plane which do not contain any images such that the observed image properties remain unchanged.
@@ 4.3 Generalized mass-sheet degeneracy (gMSD)
As discussed in the previous sub-section, in MSD, we add a constant mass density, which is a multiple of the critical density ($\Sigma_{\rm cr}$). However, exactly due to the fact that it is multiple of critical density, it can be broken if we have two lensed sources since critical density is redshift dependent. Liesenborgs et al. (2008b); Liesenborgs & De Rijcke (2012) showed that one could construct a so-called generalized mass-sheet degeneracy (gMSD), which cannot be broken even in the presence of multiple lensed sources. This is done by first rescaling $\kappa(\boldsymbol{x})$ for each lensed source mass-sheet transformation and then adding monopoles to compensate for each mass-sheet such that each of the lensed sources will only be subjected to their own MSDs. This makes the gMSD a local degeneracy according to our definition as it re-scales the time delay surface in a position-dependent manner. As shown in sub-section 4.1, adding monopoles essentially modifies the time delay surface (although not at the image position), and these changes will again be reflected in the $F(t)$ and $F(f)$ curves. Hence, we can again argue that, in wave-optics, the gMSD will be broken.
@@ 5 Summary & Conclusion
In this work, we have studied various lensing degeneracies in the wave-optics regime, primarily focusing on lensed GW signals. We have explicitly shown that monopole degeneracy breaks in the wave-optics regime since its effects are captured in the amplification factor. Based on that, we argue that any local degeneracy can be broken in the wave-optics regime since, by definition, they will modify the arrival time delay surface (although not at the image positions) as well as $F(f)$. As mentioned in 4.1, the frequency range that is used to break the monopole degeneracy is directly related to the substructure properties as it needs to affect the corresponding time delay contours.
@@ Acknowledgements
I am grateful to Prasenjit Saha for many useful discussions and comments. I also thank Jose Diego and Jenny Wagner for their useful comments. Author also thanks the reviewer for useful comments. I acknowledge support by grant 2020750 from the United States-Israel Binational Science Foundation (BSF), grant 2109066 from the United States National Science Foundation (NSF), and the Ministry of Science & Technology, Israel. This research has made use of NASA’s Astrophysics Data System Bibliographic Services.